OpenStudy (anonymous):

A cylinder’s radius is reduced to 2/5 its original size and the height is quadrupled. How has the volume of the cylinder changed? WIll fan if u answer.

OpenStudy (anonymous):

For a cylinder, \[V=\pi r^2 h\]Our new cylinder has volume\[V_1=\pi (\frac{2r}{5})^2(4h)=\frac{16}{25} \pi r^2 h\]It would seem the new cylinder has about 16/25 the volume of the original.